A body takes n times as much time to slide down a 45degree rough incline as it takes to slide down a similar but perfectly smooth 45 degree incline. the coefficient of friction between the body and the incline is?

Grade:11

1 Answers

Kenny Toijam
7 years ago
S=(1/2)gsin45 t2-------------------(1)S=(1/2)g(sin45-cos45x)(nt)2 -------------------(2) where x=coeff of frictionHere S=S or, x=1-(1/n)2